OpenStudy (anonymous):

Simplify 12k+8k+15j-4k+6j-3j

OpenStudy (anonymous):

12k+8k+15j-4k+6j-3j =20k+15j-4k+6j-3j =20k-4k+15j+6j-3j =16k+15j+6j-3j =16k+21j-3j =16k+18j